OpenStudy (anonymous):

Find the equation of a line through (4, -3) that is parallel to the line that passes through these points (-4,7) and (2,4). Write your answer in standard form (Ax + By = C). Show all your work.

OpenStudy (anonymous):

I think I know how to do it I just need some help in directing me in what to start with.

OpenStudy (texaschic101):

using the two points (-4,7) and (2,4)....find the slope

OpenStudy (texaschic101):

do you know how to do that ?

OpenStudy (texaschic101):

do you need me to show you ?

OpenStudy (texaschic101):

use the slope formula.. slope(m) = (y2 - y1) / (x2 - x1) (-4,7)...x1 = -4 and y1 = 7 (2,4)...x2 = 2 and y2 = 4 now sub slope(m) = (4 - 7) / (2 - (-4) slope(m) = -3/(2 + 4) slope(m) = -3/6 slope(m) = -1/2 a parallel line will have the same slope....so since we already know the slope(m) we need to find the y intercept (b) y = mx + b slope(m) = -1/2 (4,-3)..x = 4 and y = -3 now we sub and solve for b, the y intercept -3 = -1/2(4) + b -3 = - 2 + b -3 + 2 = b -1 = b so our parallel equation is : y = -1/2x - 1 But we need it in standard form Ax + By = C y = -1/2x - 1 --- multiply everything by 2 2y = -x - 2 -- add x to both sides x + 2y = -2 <===standard form

OpenStudy (anonymous):

Okay, thank you so much! I am going to work through it myself again and make sure I am doing all the steps right but this helps a lot.